OpenStudy (anonymous):

precalc givin medals 1. Let u = <-4, 1>, v = <-1, 6>. Find -2u + 4v. <4, 22> <4, 7> <12, -26> <10, -14> 2. Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <3, 0>, v = <0, -6> Orthogonal Neither Parallel

OpenStudy (freckles):

u is a vector the -2 in front of the vector u is a scale factor it is just a number you multiply to each component of u

OpenStudy (freckles):

from number 2 have you considered finding the dot product?

OpenStudy (freckles):

and the magnitudes of the vectors as well

OpenStudy (freckles):

\[\cos(\theta)=\frac{a \cdot b}{|a| \cdot |b|} \\ \theta=\arccos( \frac{ a \cdot b}{|a| \cdot |b|}) \\ \text{ if } \theta=90^o \text{ then } a \text{ and } b \text{ are orthogonal } \\ \text{ but to find out if they are orthogonal you only need to see if } a \cdot b=0 \\ \text{ this is much shorter than actually finding the angle between the two vectors }\] \[\text{ if } \theta=0^o \text{ or } \theta=180^o \text{ then you have that } a \text{ and } b \\ \text{ are parallel } \\ \text{ but you don't really need to find } \theta \\ \text{ you could just see if } a \cdot b=|a| \cdot |b| \text{ or if } a \cdot b =-|a| \cdot |b|\]

OpenStudy (anonymous):

@freckles orthogonal?

OpenStudy (dan815):

u and v are definately orth as one is completely in x and the onther is in y

OpenStudy (dan815):